# Amputation-imputation based generation of synthetic tabular data for ratemaking

### v1

@@ 3 Methodology
3.1 Multiple imputation by chained equations
According to van Buuren, (2018), multiple imputation by chained equations (MICE) is one of the best approaches to dealing with missing data. MICE is also known as fully conditional specification, sequential regression multivariate imputation, or simply chained equations.
MICE commonly operates under the assumption that data is missing at random, which means that the probability of a value being missing depends only on observed values and not on unobserved values. Avoiding explicit assumptions on the joint distribution of the data, MICE works by specifying a set of conditional distributions for each variable with missing data, given the other variables in the data. The algorithm then iteratively imputes the missing values by drawing from these conditional distributions. This process of iterative drawing can be seen as a Gibbs sampler that will converge to the draws from the theoretical joint distribution of data, should this joint distribution exist (cf. Raghunathan et al., (2001)). In practice, the assumption about the existence of the joint distribution is often implicitly made.
Borrowing notation from Drechsler, (2011), we summarize the MICE algorithm:
@@ 3.1 Multiple imputation by chained equations
    Iterate Steps 2-4 for several cycles to update imputations until the convergence or the maximal number $n_{MI}$ of MICE iterations is reached.
@@
[…]
@@
Once multiple imputed datasets have been created, inference can be done regarding a quantity of interest (e.g., mean of a variable, correlation coefficient between two variables). So-called Rubin’s rules provide statistical foundation for such inference. For a scheme that illustrates how MICE works, see Figure 1. For more information, see, e.g., Section 5.1 in Drechsler, (2011).
@@
As mentioned earlier, the MICE framework can be used to generate synthetic datasets. For simplicity, we assume that the training data has no missing values, otherwise, a 2-step approach is needed, where before the generation (in the second step) in the first step, missing values are imputed. The idea is to iteratively select parts of the original data and then replace them with synthetic values drawn from the posterior distribution learned by MICE.
@@
There are different ways to sequentially choose chunks of data for replacement (imputation). One way is to choose a full column in each iteration, cf. Volker and Vink, (2021). This case has two peculiarities. First, the imputation model is trained on the observed data, which is then replaced by the model prediction. Second, the MICE algorithm converges immediately and each draw will be a direct draw from the posterior distribution, because a so-called monotone missingness pattern is satisfied for column-wise replacement/imputation (see Section 3.1.2 in Drechsler, (2011)). Another approach, which in general does not satisfy the monotone missingness pattern, is to randomly select in each amputation-imputation step cells whose values must be replaced. The values in these cells are deleted, interpreted as missing, and the MICE algorithm is executed, requiring more than one iteration to converge.
@@
In our case studies, we test several ways of iteratively amputating data and then applying MICE to impute it. One of them is inspired by the tabulator approach of Neves et al., (2022), where the authors suggest to amputate data in chunks of approximately the same size and to impute deleted data using a GAIN. Instead of a GAIN, we will use MICE as the imputation engine.
@@
Similarly to the MICE-based imputation, there are Rubin rules for analyzing (e.g., computing a point estimate, an estimate for the variance) unknown statistical quantities from fully synthetic datasets, cf. Reiter, (2003) and Section 6.1 in Drechsler, (2011).
@@
In our case studies, we use random forests (RFs) as the imputation model. This imputation model and many others are implemented in the R-package mice, cf. van Buuren and Groothuis-Oudshoorn, (2011). RFs were introduced in Breiman, (2001). It is a non-parametric model that can capture complex relationships among variables without the need to specify a particular regression model. This makes RFs well-suited for imputing missing data in datasets with non-linear interactions among variables. Another important advantage of RFs is that they can often be used out-of-the-box and can be used rather easily to handle heterogeneous data. For more information on RFs in actuarial context, see, e.g., Chapter 4 in Denuit and Trufin, (2020) or Chapter 6 in Wuthrich et al., (2025).
@@ 3.2 Conditional Tabular GAN
In this subsection, which is based on Kuo, (2020) and Sections $2$ and $3$ of Cote et al., (2020), we describe how a GAN works and explain how a CTGAN extends it.
@@
Proposed in Xu et al., (2019), CTGAN builds upon the GAN framework by incorporating several innovations to address the challenges of generating synthetic tabular data: mode-specific normalization, conditional generation of data, and handling of mixed data. Next we summarize an exemplary architecture of CTGAN, as in Cote et al., (2020).
As per Kuo, (2020), a CTGAN generates individual records by randomly selecting a variable (e.g., fuel type) and then choosing a value (e.g., diesel) based on the actual data frequency. The algorithm finds a matching training data row for this value and generates the remaining variables conditioned on the initial selection. Both the generated and true records are evaluated by the discriminator. Figure 3 in Cote et al., (2020) schematically illustrates this process and Figure 2 in Cote et al., (2020) visually exemplifies the architecture of the generator and discriminator. The discriminator (blue) and the generator (orange) each have two fully-connected layers to capture relationships between columns. The generator, which supports categorical variables, employs skip connections; see Xu et al., (2019) for extensions to continuous variables. CTGAN also uses the PacGAN framework in the discriminator to mitigate a so-called mode collapse. This peculiarity means that the model focuses on a small subset of the real data and, thus, has a bias towards the generation of samples that are not sufficiently diverse.
Although a CTGAN tends to be substantially better than a GAN for generating high-cardinality categorical variables, this type of variables can be still challenging for a CTGAN, depending on the characteristics of the real data such as the degree of cardinality and the frequency of each category. Aiming to improve the performance of a CTGAN for high-cardinality categorical variables, we try encoding them via non-linear autoencoders (AEs) before training a CTGAN.
@@ 3.3 Autoencoders
The original aim of AEs is learning a compressed representation of data (encoding) and then reconstructing the original input (decoding). Structurally, an AE can be thought of as a pair of connected feedforward NNs. So principal components analysis approach can be interpreted as a linear AE, i.e., the activation functions of both NNs are linear. Hinton and Salakhutdinov, (2006) demonstrated the power of non-linear AEs for learning better, lower-dimensional representations than linear AEs. In our paper, we use two different types of AEs due to their different application – deterministic AEs and variational AEs.
Our application of deterministic AEs is the transformation of high-cardinality categorical variables to low-dimensional numerical vectors (instead of the one-hot encoding) before fitting them to GAN-based models for data generation. The benefits of using such AEs for training NNs in insurance predictive modeling has been studied in Delong and Kozak, (2023), which is our main reference for this application of non-linear AEs.
@@
An AE is trained in an unsupervised manner. If an AE achieves a small reconstruction error, it is said that the encoder $\varphi$ extracts the most important information from a multi-dimensional vector of features $\bm{a}$. In our application, $\bm{a}$ will be a vector of features coming from a one-hot encoded categorical covariate, i.e., $\bm{a}=\bm{x}_{j}^{cat}$ with $\bm{x}_{j}^{cat}$ defined as per (2.1), $j\in\left\{1,\dots,n_{CC}\right\}$. For each categorical variable $\bm{x}_{j}^{cat}$, a separate AE is trained to reduce the dimensionality $p_{j}=m_{j}$ of $\bm{x}_{j}^{cat}$ to $l_{j}<p_{j}$, $j\in\left\{1,\dots,n_{CC}\right\}$. Afterwards, this representation $\overline{\bm{x}}_{j}^{cat}=\varphi\left({\bm{x}}_{j}^{cat}\right)$, instead of ${\bm{x}}_{j}^{cat}$, is used as input for CTGAN to generate data. The approach of using a separate AE for each categorical variable is called Separate AEs in Delong and Kozak, (2023). We use the same architecture of separate AEs, as proposed by Delong and Kozak, (2023). In particular, the reconstruction function $\pi$ is an NN with one hidden layer. The output-layer neurons have softmax activation functions and return the probabilities that the reconstructed feature takes a particular category. The category with the highest predicted probability is the one predicted for the reconstructed feature. $L$ in (3.2) is a cross-entropy loss function.
The approach described above is applied to all categorical features in the dataset. An example of the architecture of a neural network used in this paper to build the autoencoder of type Separate AEs for categorical features (with 2 and 3 categories) is presented in Figure 2 in Delong and Kozak, (2023). See their Section 3.2 for further details on Separate AEs.
Instead of using a separate AE for each categorical variable, we could have one joint AE for all categorical variables simultaneously. This approach is called Joint AE in Delong and Kozak, (2023). According to Figure 4 there, for a large number of epochs (e.g., $100$ for freMTPL2freq data) and a sufficiently high dimension of the representation of all categorical features (e.g., $15$ for freMTPL2freq data), the reconstruction error power measured with the cosine similarity is very similar for Separate AEs and Joint AE.
The second type of AEs used in our paper is variational autoencoders (VAEs). They were introduced by Kingma and Welling, (2013) and can be seen as a Bayesian extension to deterministic AEs. VAEs learn a probabilistic representation of data points in a latent space, which is why these models can be used for synthetic data generation in addition to representation learning. In our case studies, we also consider the VAE-based approach proposed by Jamotton and Hainaut, (2024), who also tested their methodology on a data set with about $62000$ observations. For an overview of VAEs and details about their methodology, we refer the reader to their article.
@@ 4 Case studies
This section is dedicated to our comparative study, which is based on freMTPL2freq dataset. In Subsection 4.1, we provide the details about our method training pipelines, the mechanics of data augmentation, and the evaluation pipelines for generation methods. In Subsection 4.2, we specify all the methods tested and state explicit formulas for modeling the structural dependence between observed covariates and simulated claim counts that represent “ground truth” in tested scenarios. The results are discussed in Subsection 4.3.
@@ 4.1 General setting and evaluation metrics
Our aim with this case study is to answer the research questions stated at the end of Section 2 and, in doing so, utilize pre-existing implementations of synthetic data generation methods with minimal user interventions for data preparation or method training. This aspect is significant in practice since, in our experience from the industry, actuaries seldom have sufficient time to implement and fine-tune their specific algorithms and methods. Since we want to identify best “off-the-shelf” data generation approach, we do not do any pre-processing of variables when preparing data in this section. Any processing is done by the generation method implementation. We execute our case study on the high-performance computation cluster at the University of Tartu and can provide our code upon request.
In our experiments, we use freMTPL2freq dataset to give us a baseline of what a realistic dataset might look like. It has $678013$ observations with $9$ explanatory variables ($5$ categorical and $4$ numeric), $1$ exposure variable and $1$ response variable (the number of claims). This dataset has been used in multiple actuarial data science articles, including Kuo, (2020), where it is used for training a CTGAN to generate synthetic data.
@@
In the second formula we use the previously defined ${f}_{lin}$ and additionally assume a set of interactions among covariates, namely ${f}_{int}(\bm{x})={f}_{lin}(\bm{x})+\sum_{\{j,k\}\in\mathcal{I}}\beta_{j;k}x_{j}x_{k}$, where the interaction part is defined as in Equation (2.3) and $\mathcal{I}$ is the set of index pairs for the interactions of interest. Although we use double indices for notational simplicity, one can construct a simple mapping from two indices to single indices. The values ${f}_{lin}(\bm{x}_{i})$ and ${f}_{int}(\bm{x}_{i})$ are then used as the means of the respective Poisson distributions to simulate the number of claims $y_{i}$, $i\in\left\{1,\dots,n_{ROD}\right\}$.
So, in our experiments, we have two datasets, each of which contains simulated number claims for which we know their “true” structural dependence on original covariates. In the sequel, we refer to this artificially generated number of claims as the response variable and use the $90\%-10\%$ training-test split, i.e., the number of rows in the train set is $n_{RTR}=\lfloor n_{ROD}\cdot 0.9\rfloor$.
Let now $\mathcal{A}=\left\{A_{1},\dots,A_{n_{A}}\right\}$ be the set of algorithms (approaches) for generation of synthetic data, $n_{A}\in\mathbb{N}$ is the number of algorithms for synthetic data generation. For each approach $A$, we denote by $GM_{A}$ as the generation method for this algorithm.
@@
Each generation method is used to generate $n_{E}$ (referred to as the number of experiments) synthetic datasets, where the number of rows for a single synthetic dataset was the same as the training data. Denote a generated synthetic dataset as $\mathcal{S}^{A;k}$, where $k\in\{1,\dots,n_{E}\}$ and $A\in\mathcal{A}$.
@@
If $x_{j}^{orig}$ is numeric, then it is discretized into $n_{bin_{j}}$ bins, where bin cut-off values are selected based on the quantiles of the numeric variable values in the original training data. The quantiles are used to ensure that the proportion of observations in each bin is the same or very close. Note that in the case when one value appears in several quantiles (e.g., one value is very prevalent in the data), then it has a bin all on its own. For each bin, then, a binary variable (feature) is added, and the vector of these features is denoted by $\bm{x}_{j}^{cat}$. Additionally, we denote by $\mathcal{I}_{j}$ the set of indices for preprocessed variables corresponding to $\bm{x}_{j}^{cat}$ .
@@
For the evaluation of model metrics, the underlying models need to be trained. We focus on using GLMs as the predictive model type. For each dataset $\mathcal{S}^{A;k;\bm{s}}$, two predictive GLM models are derived. One model, where the formula used to build the model is the same as the corresponding true model formula (i.e. linear formula or interactions formula), and one which is derived using a bidirectional step-wise variable selection using $AIC$. Note that for interaction formula, the scope for the variable selection included variables explicitly defining the interactions between the different variables (since this was the easiest way to include interactions in the step-wise search). The values related to the trained GLM models are denoted as follows:
@@
  • $\widehat{\bm{\beta}}^{A;k;\bm{s}}=\left(\widehat{\beta}_{0}^{A;k;\bm{s}},\widehat{\beta}_{1}^{A;k;\bm{s}},\dots,\widehat{\beta}_{d}^{A;k;\bm{s}}\right)^{\top}$ – the vector of estimated model coefficients based on pre-specified true model structure, using the GLM trained on trained on the dataset specified by the data structure parameter $s$.
@@
As these values can be obtained for every trained GLM model, we use two metrics to assess the effect of synthetic data on estimated values of model coefficients.

### v2

@@ 3 Methodology
3.1 Multivariate imputation by chained equations
According to van Buuren (2018), multivariate imputation by chained equations (MICE) is one of the best approaches to dealing with missing data. MICE is also known as fully conditional specification, sequential regression multivariate imputation, or simply chained equations.
MICE commonly operates under the assumption that data is missing at random, which means that the probability of a value being missing depends only on observed values and not on unobserved values. Avoiding explicit assumptions on the joint distribution of the data, MICE works by specifying a set of conditional distributions for each variable with missing data, given the other variables in the data. The algorithm then iteratively imputes the missing values by drawing from these conditional distributions. This process of iterative drawing can be seen as a Gibbs sampler that will converge to the draws from the theoretical joint distribution of data, should this joint distribution exist (cf. Raghunathan et al. (2001)). In practice, the assumption about the existence of the joint distribution is often implicitly made.
Borrowing notation from Drechsler (2011), we summarize the MICE algorithm:
@@ 3.1 Multivariate imputation by chained equations
    Iterate Steps 2-5 for several cycles to update imputations until the convergence or the maximal number $n_{MI}$ of MICE iterations is reached.
@@
Figure 1: Schematic representation of MICE. $\bm{Y}^{(t-1)}_{-j}:=\left(Y^{(t)}_{1},\dots,Y^{(t)}_{j-1},Y^{(t-1)}_{j+1},\dots,Y^{(t-1)}_{p}\right)$, $Y^{(t)}_{j,mis}$ denotes the elements of the column corresponding to $Y^{(t)}_{j}$ that must be imputed ($\Bigl$(i.e., missing in the column corresponding to $Y^{(0)}_{j}\Bigr)$, $Y^{(t)}_{j,obs}$ denotes the elements in the respective column that were originally observed in $Y^{(0)}_{j}$, $f_{j}$ is an imputation model (e.g., a GLM, a random forest, etc.), $\bm{\theta}^{(t)}_{j}$ is the vector of parameters of $f_{j}$ in MICE iteration $t$.
@@
Once multiple imputed datasets have been created, inference can be done regarding a quantity of interest (e.g., the mean of a variable, the correlation coefficient between two variables). The so-called Rubin’s rules provide a statistical foundation for such inference. For a scheme that illustrates how MICE works, see Figure 1. For more information, see, e.g., Section 5.1 in Drechsler (2011).
@@
As mentioned earlier, the MICE framework can be used to generate synthetic datasets. For simplicity, we assume that the training data has no missing values; otherwise, a 2-step approach is needed, where missing values are imputed in the first step before generation (second step). The idea is to iteratively select parts of the original data and then replace them with synthetic values drawn from the posterior distribution learned by MICE.
@@
There are different ways to sequentially choose chunks of data for replacement (imputation). One way is to choose a full column in each iteration, cf. Volker and Vink (2021). This case has two peculiarities. First, the imputation model is trained on the observed data, which is then replaced by the model prediction. Second, the MICE algorithm converges immediately, and each draw will be a direct draw from the posterior distribution, because a so-called monotone missingness pattern is satisfied for column-wise replacement/imputation (see Section 3.1.2 in Drechsler (2011)). Another approach, which in general does not satisfy the monotone missingness pattern, is at each amputation-imputation step to randomly select cells whose values must be replaced. The values in these cells are deleted, and treated as missing, and the MICE algorithm is executed, requiring more than one iteration to converge.
@@
In our case studies, we test several ways to iteratively set a portion of data to missing and then apply MICE to impute. One of them is inspired by the tabulator approach of Neves et al. (2022), where the authors suggest setting data to missing in chunks of approximately equal size and to impute missing data using a GAIN. Instead of a GAIN, we will use MICE as the imputation engine.
@@
Similarly to the MICE-based imputation, there are Rubin rules for analyzing (e.g., computing a point estimate, an estimate for the variance) unknown statistical quantities from fully synthetic datasets, cf. Reiter (2003) and Section 6.1 in Drechsler (2011).
@@
In our case studies, we use random forests (RFs) as the imputation model. This imputation model and many others are implemented in the R-package mice, cf. van Buuren and Groothuis-Oudshoorn (2011). RFs were introduced in Breiman (2001). It is a non-parametric model that can capture complex relationships among variables without the need to specify a particular regression model. This makes RFs well-suited for imputing missing data in datasets with non-linear interactions among variables. Another important advantage of RFs is that they can often be used out-of-the-box and can be used rather easily to handle heterogeneous data. For more information on RFs in actuarial context, see, e.g., Chapter 4 in Denuit and Trufin (2020) or Chapter 6 in Wuthrich et al. (2025).
@@ 3.2 Conditional Tabular GAN
In this subsection, which is based on Kuo (2020) and Sections $2$ and $3$ of Côté et al. (2025), we describe how a GAN works and explain how a CTGAN extends it.
@@
Proposed in Xu et al. (2019), CTGAN builds upon the GAN framework by incorporating several innovations to address the challenges of generating synthetic tabular data: mode-specific normalization, conditional generation of data, and handling of mixed data. Next we summarize an exemplary architecture of CTGAN, as in Côté et al. (2025).
As per Kuo (2020), a CTGAN generates individual records by randomly selecting a variable (e.g., fuel type) and then choosing a value (e.g., diesel) based on the actual data frequency. The algorithm finds a matching training data row for this value and generates the remaining variables conditioned on the initial selection. Both the generated and true records are evaluated by the discriminator. Figure 3 in Côté et al. (2025) schematically illustrates this process and Figure 2 in Côté et al. (2025) visually exemplifies the architecture of the generator and discriminator. The discriminator (blue) and the generator (orange) each have two fully-connected layers to capture relationships between columns. The generator, which supports categorical variables, employs skip connections; see Xu et al. (2019) for extensions to continuous variables. CTGAN also uses the Packed Generative Adversarial Network (PacGAN) framework in the discriminator to mitigate a so-called mode collapse. This peculiarity means that the model focuses on a small subset of the real data and, thus, has a bias towards the generation of samples that are not sufficiently diverse.
Although a CTGAN tends to be substantially better than a GAN for generating high-cardinality categorical variables, this type of variables can still be challenging for a CTGAN, depending on the characteristics of the real data, such as the degree of cardinality and the frequency of each category. Aiming to improve the performance of a CTGAN for high-cardinality categorical variables, we try encoding them via non-linear autoencoders (AEs) before training a CTGAN.
@@ 3.3 Autoencoders
The original aim of AEs is learning a compressed representation of data (encoding) and then reconstructing the original input (decoding). Structurally, an AE can be thought of as a pair of connected feedforward NNs. So principal components analysis approach can be interpreted as a linear AE, i.e., the activation functions of both NNs are linear. Hinton and Salakhutdinov (2006) demonstrated the power of non-linear AEs for learning better, lower-dimensional representations than linear AEs. In our paper, we use two different types of AEs due to their different application – deterministic AEs and variational AEs.
Our application of deterministic AEs is the transformation of high-cardinality categorical variables to low-dimensional numerical vectors (instead of the one-hot encoding) before fitting them to GAN-based models for data generation. The benefits of using such AEs for training NNs in insurance predictive modeling have been studied in Delong and Kozak (2023), which is our main reference for this application of non-linear AEs.
@@
An AE is trained in an unsupervised manner. If an AE achieves a small reconstruction error, it is said that the encoder $\varphi$ extracts the most important information from a multi-dimensional vector of features $\bm{a}$. In our application, $\bm{a}$ will be a vector of features coming from a one-hot encoded categorical covariate, i.e., $\bm{a}=\bm{x}_{j}^{cat}$ with $\bm{x}_{j}^{cat}$ defined as per (2.1), $j\in\left\{1,\dots,n_{CC}\right\}$. For each categorical variable $\bm{x}_{j}^{cat}$, a separate AE is trained to reduce the dimensionality $p_{j}=m_{j}$ of $\bm{x}_{j}^{cat}$ to $l_{j}<p_{j}$, $j\in\left\{1,\dots,n_{CC}\right\}$. Afterwards, this representation $\overline{\bm{x}}_{j}^{cat}=\varphi\left({\bm{x}}_{j}^{cat}\right)$, instead of ${\bm{x}}_{j}^{cat}$, is used as input for CTGAN to generate data. The approach of using a separate AE for each categorical variable is called Separate AEs in Delong and Kozak (2023). We use the same architecture of separate AEs, as proposed by Delong and Kozak (2023). In particular, the reconstruction function $\pi$ is an NN with one hidden layer. The output-layer neurons have softmax activation functions and return the probabilities that the reconstructed feature takes a particular category. The category with the highest predicted probability is the one predicted for the reconstructed feature. $L$ in (3.2) is a cross-entropy loss function.
The approach described above is applied to all categorical features in the dataset. An example of the architecture of a neural network used in this paper to build the autoencoder of type Separate AEs for categorical features (with 2 and 3 categories) is presented in Figure 2 in Delong and Kozak (2023). See their Section 3.2 for further details on Separate AEs.
Instead of using a separate AE for each categorical variable, we could have one joint AE for all categorical variables simultaneously. This approach is called Joint AE in Delong and Kozak (2023). According to Figure 4 there, for a large number of epochs (e.g., $100$ for the freMTPL2freq data) and a sufficiently high dimension of the representation of all categorical features (e.g., $15$ for the freMTPL2freq data), the reconstruction error power measured with the cosine similarity is very similar for Separate AEs and Joint AE.
The second type of AEs used in our paper is variational autoencoders (VAEs). They were introduced by Kingma and Welling (2013) and can be seen as a Bayesian extension to deterministic AEs. VAEs learn a probabilistic representation of data points in a latent space, which is why these models can be used for synthetic data generation in addition to representation learning. In our case studies, we also consider the VAE-based approach proposed by Jamotton and Hainaut (2024), who also tested their methodology on a dataset with about $62000$ observations. For an overview of VAEs and details about their methodology, we refer the reader to their article.
@@ 4 Case studies
This section is dedicated to our comparative study, which is based on the freMTPL2freq dataset. In Subsection 4.1, we provide the details about our method training pipelines, the mechanics of data augmentation, and the evaluation pipelines for generation methods. In Subsection 4.2, we specify all the methods tested and state explicit formulas for modeling the structural dependence between observed covariates and simulated claim counts that represent “ground truth” in tested scenarios. The results are discussed in Subsection 4.3.
@@ 4.1 General setting and evaluation metrics
Our aim with this case study is to answer the research questions stated at the end of Section 2 and, in doing so, utilize pre-existing implementations of synthetic data generation methods with minimal user interventions for data preparation or method training. This aspect is significant in practice since, in our experience from the industry, actuaries seldom have sufficient time to implement and fine-tune their specific algorithms and methods. Since we want to identify the best “off-the-shelf” data generation approach, we do not do any pre-processing of variables when preparing data in this section. Any processing is done by the generation method implementation. We execute our case study on the high-performance computation cluster at the University of Tartu and can provide our code upon request.
In our experiments, we use the freMTPL2freq dataset to give us a baseline of what a realistic dataset might look like. It has $678013$ observations with $9$ explanatory variables ($5$ categorical and $4$ numeric), $1$ exposure variable and $1$ response variable (the number of claims). This dataset has been used in multiple actuarial data science articles, including Kuo (2020), where it is used for training a CTGAN to generate synthetic data. The freMTPL2freq dataset is the de facto benchmark in modern actuarial data science, and introducing non-standard datasets would reduce comparability with existing literature.
@@
In the second formula we use the previously defined ${f}_{lin}$ and additionally assume a set of interactions among covariates, namely ${f}_{int}(\bm{x})={f}_{lin}(\bm{x})+\sum_{\{j,k\}\in\mathcal{I}}\beta_{j;k}x_{j}x_{k}$, where the interaction part is defined as in Equation (2.3) and $\mathcal{I}$ is the set of index pairs for the interactions of interest. Although we use double indices for notational clarity, one can construct a simple mapping from double indices to single indices. The values ${f}_{lin}(\bm{x}_{i})$ and ${f}_{int}(\bm{x}_{i})$ are then used as the means of the respective Poisson distributions to simulate the number of claims $y_{i}$, $i\in\left\{1,\dots,n_{ROD}\right\}$.
So, in our experiments, we have two datasets, each of which contains simulated number claims for which we know their “true” structural dependence on original covariates. In the sequel, we refer to this artificially generated number of claims as the response variable and use the $90\%/10\%$ training/test split, i.e., the number of rows in the train set is $n_{RTR}=\lfloor n_{ROD}\cdot 0.9\rfloor$.
Let now $\mathcal{A}=\left\{A_{1},\dots,A_{n_{A}}\right\}$ be the set of algorithms (approaches) for generation of synthetic data, $n_{A}\in\mathbb{N}$ is the number of algorithms for synthetic data generation. For each approach $A$, we denote the generation method for this algorithm by $GM_{A}$.
@@
Each generation method is used to generate $n_{E}$ (referred to as the number of experiments) synthetic datasets, where the number of rows for a single synthetic dataset is the same as in the training data. Denote a generated synthetic dataset as $\mathcal{S}^{A;k}$, where $k\in\{1,\dots,n_{E}\}$ and $A\in\mathcal{A}$.
@@
If $x_{j}^{orig}$ is numeric, then it is discretized into $n_{bin_{j}}$ bins, where bin cut-off values are selected based on the quantiles of the numeric variable values in the original training data. The quantiles are used to ensure that the proportion of observations in each bin is the same or very close. Note that if a value appears in several quantiles (e.g., it is very prevalent in the data), it has its own bin. For each bin, then, a binary variable (feature) is added, and the vector of these features is denoted by $\bm{x}_{j}^{cat}$. Additionally, we denote by $\mathcal{I}_{j}$ the set of indices for preprocessed variables corresponding to $\bm{x}_{j}^{cat}$ .
@@
For the evaluation of model metrics, the underlying models need to be trained. We focus on using GLMs as the predictive model type. For each dataset $\mathcal{S}^{A;k;\bm{s}}$, two predictive GLM models are derived. One model, where the formula used to build the model is the same as the corresponding true model formula (i.e., linear formula or interaction formula), and one which is derived using a bidirectional step-wise variable selection using $AIC$. Note that for the interaction formula, variable selection included explicitly defined columns for each of the interactions between the different variables, as this was the easiest way to include interactions in the step-wise search. The values related to the trained GLM models are denoted as follows:
@@
  • $\widehat{\bm{\beta}}^{A;k;\bm{s}}=\left(\widehat{\beta}_{0}^{A;k;\bm{s}},\widehat{\beta}_{1}^{A;k;\bm{s}},\dots,\widehat{\beta}_{d}^{A;k;\bm{s}}\right)^{\top}$ – the vector of estimated model coefficients based on pre-specified true model structure, using the GLM trained on the dataset specified by the data structure parameter $s$.
@@
As these values can be obtained for every trained GLM model, we use two metrics to assess the effect of synthetic data on the estimated values of model coefficients.